A 15.0% by mass solution of glucose (C 6H 12O 6) has a density of 1.10 g/mL. What is the molarity?
Answer #1

Molar mass of glucose (C6H12O6) = 6x12 + 12x1 + 6x16 =180 g /mol

Number of moles of glucose = 15/180 = 0.0833 moles

Molarity = Number of moles of solute / Volume of solution in liter
Density = 1.10 g/ml
Volume = mass / density
= 100 g / 1.10 g/ ml = 90.90 ml = 0.09090 L
Molarity = 0.0833/ 0.09090
= 0.916 M